OpenStudy (anonymous):

Determine the rate of change: -2y-16x=41

OpenStudy (anonymous):

well first you want to put it into point slope form and simplify it so you would subtract 41 from both sides giving you (-2y-16x-41=0) and then add 2y to both sides giving you (-16x-41=2y) Now just simplify it by dividing everything by 2 giving you -8x-20.5(20 1/2)=y and -8 is the rate of change because the rate of change is just technically your slope.